OpenStudy (awesomemiss):

Convert 5 over 8 to a decimal and tell whether it terminates or repeats. 5.8, a number that terminates 5.8 with a repeating bar over 8., a number that repeats 0.625 with a repeating bar over 625., a number that repeats 0.625, a number that terminates

OpenStudy (awesomemiss):

@DallasCowboys88 @ganeshie8 @ILovePuppiesLol @welshfella @mathmate please help me

OpenStudy (quickstudent):

The easiest way to convert a fraction to a decimal is to divide the numerator by the denominator. This can be done easily on a calculator:)

OpenStudy (quickstudent):

So for this question, simply divide 5 by 8:) Let's see your answer:)

OpenStudy (awesomemiss):

i know what the decimal is its 0.625 but im having trouble seeing if it terminates or repeats

OpenStudy (awesomemiss):

i think its D

OpenStudy (quickstudent):

Repeats means that the same number goes on forever. If it just stops at 0.625, then it terminates:)

OpenStudy (awesomemiss):

ok thanks so its D

OpenStudy (quickstudent):

Correct:)

OpenStudy (mathmate):

@awesomemiss You can check if the decimal repeats without actually doing a division! If the denominator is made up of powers of 2 and/or 5, the fraction does not repeat. If it contains any factor other than 2 or 5, the resulting fraction repeats. Example: 7/50 does not repeat, because 50=2*5*5, i.e. product of powers of 2 and 5. 39/384 repeats, because 384=128*3=2^7*3. Since we have a factor of 3 (not 2, not 5), the fraction repeats.